OpenStudy (anonymous):

is 300 to 450 the same as 250 to 300

OpenStudy (anonymous):

no

OpenStudy (anonymous):

No, it's not equal. Source: http://www.wolframalpha.com/input/?i=300^450+%3D+250^300 But if you don't have a super-calculator like wolfram alpha, you think the following, You know that: \[a < b \Rightarrow a ^{300} < b^{300} \Rightarrow a^{450} < b^{300}\] Consider a= 250 and b=300

OpenStudy (jdoe0001):

\(\bf \cfrac{300}{450}= \cfrac{250}{300}\qquad ?\) what do you think?

OpenStudy (anonymous):

jdoe0001, that's also false. Source: http://www.wolframalpha.com/input/?i=%28300%2F450%29+%3D+%28250%2F300%29 You can check that by doing the following: \[\frac{ 300 }{ 450 } \neq \frac{250}{300}\] \[\frac{300}{450} \times \frac{300}{300} \neq \frac{250}{300} \times \frac{450}{450} \] \[\frac{90000}{135000} \neq \frac{112500}{135000}\]